OpenStudy (anonymous):

Find an equation for the line that is perpendicular to y=4x-5. and passes through (5, -2)

OpenStudy (anonymous):

the slope of the line = -1/m where m is the slope of your line use the general form y-y1 =m(x-x1) where x1 y1 is the point on the line and m is the slope you have just found

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

do you know the slope of the line y = 4x - 5?

OpenStudy (anonymous):

4x

OpenStudy (anonymous):

4 is the slope - the coefficient of x ( ie number that multiplies x)

OpenStudy (anonymous):

so slope of perpendicular line = -1/4

OpenStudy (anonymous):

now plug in -1/4 and (5,-2) into the general equation and you've got it

OpenStudy (anonymous):

-1/4x-3?

OpenStudy (anonymous):

where is y?

OpenStudy (anonymous):

y-y1 = (-1/4) (x-x1) is the equation you need to use what values go in for x1 and x1?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

yes x1 = 5 and i meant x1 and y1 - what is the value for y1?

OpenStudy (anonymous):

we are given the point (5,-2) which is on the line so (x1,y1) =( 5,?)

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

right so plug x1=5 and y1 = -2 into the equation and we have the answer

OpenStudy (anonymous):

y - 5 = (-1/4)(x - (-2))

OpenStudy (anonymous):

this can be simplified further

OpenStudy (anonymous):

thats where i got stuck,